Use the normal distribution to the right to answer the questions. Standardized Test Composite Scores (a) What percent of the scores are less than 19? (b) Out of 1500 randomly selected scores, about how many would be expected to be greater than 21? H20.9 ơ 5.8 Score (a) The percent of scores that are less than 19 is「% (Round to two decimal places as needed.) (b) Aboutscores would be expected to be greater than 21. Round to the nearest whole number as needed.)
Answer #1

Population mean, u = 20.9 Population standard deviation, o = 5.6 (a) The percent of scores are less than 19 is, X-u 19-20.9 P

(b) The probability that scores that are greater than 21 is, P(x >21)=1- P(X 521) = 1- P(Z <0.02) = 1-0.5080 = 0.492 Sample s